OpenStudy (yolo729):

how do i solve this 2^2 x 3^24

OpenStudy (mathmale):

Actually, yolo, you don't "solve this;" you evaluate it (find the value of the expression). PEMDAS *Order of operations rules" stipulate that you must do exponentiation before multiplication or division. So, your problem here breaks down as follows: 2^2 x 3^24 becomes [ 2^2]*[3^24]. Do the work inside the brackets [

OpenStudy (mathmale):

first. Then multiply. Please note that "x" in algebra is a variable; it is not for indicating multiplication.

OpenStudy (mathmale):

vague: 2^2 x 3^24 clearer: (2^2)*(3^24)

OpenStudy (mathmale):

2^2 should be relatively easy. Your result?

OpenStudy (yolo729):

2^2 = 4 3^24 = 282,429,536,481

OpenStudy (yolo729):

now do i just add

OpenStudy (yolo729):

do i add 4 + 282,429,536,481

OpenStudy (mathmale):

We were talking about multiplication, remember? You do need to add exponents, but need to say so. Basically, this is a MULT. problem. 2^2 * 3^24 = 4 * 3^24

OpenStudy (mathmale):

So, no, you do not add those two numbers together. You started with 2^2 * 3^24. First you have an exponential expression to the base 2; then you have another one, this time to the base 3. If the base is 2, the powers of 2 are always even. If the base is 3, then the powers of 3 are always odd: 3, 9, 27, 81, and so on. So, you cannot combine 4 and 3^24. All you can do is to MULTIPLY these together

OpenStudy (yolo729):

OK thank you

OpenStudy (phi):

are you sure that is the question?? because 3^24 is a huge number, and we need a calculator to figure it out. There are no "tricks" to do this (which makes me think there is a typo)

OpenStudy (mathmate):

@phi Agree! The result is of the order 10^12, which is even beyond the precision of most calculators!